OpenStudy (anonymous):

a wheelchair ramp has an angle elevation of 24 degrees. if the ramp reaches the top of a 30 inch high porch, how long is the ramp?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sin(24)=\frac{30}{x}\iff x=\frac{30}{\sin(24)}\] and a calculator

OpenStudy (anonymous):

i dont understand why x= 30/sin(24)

OpenStudy (anonymous):

solve the first equation for \(x\) and that is what you get

OpenStudy (anonymous):

x=1.25?

OpenStudy (jdoe0001):

1.25? think about it. if the porch is 30 inches high and the ramp is 1.25 long, it'd look like  so anyoneo on a wheelchair will surely fall over after the 1.25 inch length

OpenStudy (jdoe0001):

won't be really a helpful ramp, actually the guy on the wheelchair will not longer need the wheelchair, will be in bed from the fall injuries anyway

OpenStudy (anonymous):

i dont really understand how to solve for x then

OpenStudy (jdoe0001):

say \(\bf 2=\cfrac{3}{x}\) what would be the value of "x" there?

OpenStudy (anonymous):

x=3/2

OpenStudy (jdoe0001):

ahh, you see... you switch them about, well what about \(\bf sin(24^o)=\cfrac{30}{x}\)

OpenStudy (jdoe0001):

you'd do the same thing

OpenStudy (anonymous):

okay so sin(x)=30/24... then what?

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 2}}=\cfrac{3}{{\color{red}{ x}}}\implies {\color{red}{ x}}=\cfrac{3}{{\color{blue}{ 2}}}\\ \quad \\ \quad \\ {\color{blue}{ sin(24^o)}}=\cfrac{30}{{\color{red}{ x}}}\implies {\color{red}{ x}}=\cfrac{30}{{\color{blue}{ \square ?}}} \)

OpenStudy (anonymous):

\[\frac{ 30 }{\sin \left( 24 \right) } \]

OpenStudy (jdoe0001):

yeap so, that's how long the ramp is notice that satellite73 had shown that -> \(\sin(24)=\cfrac{30}{x}\iff x=\cfrac{30}{\sin(24)}\)

OpenStudy (anonymous):

@satellite73 lol loved the, and a calculator" comment.

OpenStudy (anonymous):

yeah but that still doesn't say what sin is

OpenStudy (jdoe0001):

sine is a trigonometric function

OpenStudy (jdoe0001):

check your calculator, it'd have a button shown as [sin]

OpenStudy (jdoe0001):

30 [\(\div\)] 24 [sin]

OpenStudy (anonymous):

well yes i know, but the answers are multiple choice: 97.38 32.84 12.20 73.76 none of which are a long equation

OpenStudy (jdoe0001):

well.. what did you get?

OpenStudy (anonymous):

i dont know. because if i divide 30 by 24sin i get 1.25sine

OpenStudy (anonymous):

which is way off, which is why im confused

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

ok... get the sine of 24 first stick it in the calculator's memory [M] or [M+] button usually then type in "30" then "\(\div\)" then press the Memory Recall, usually an [MR] button then enter

OpenStudy (anonymous):

30/sin(24) = 73.75780007

OpenStudy (anonymous):

ohhh, i using my calculator incorrectly.. haha sorry. thank you so much for your help. 73.76

OpenStudy (jdoe0001):

yw